OpenStudy (anonymous):

Steve has these scores on the exams for this semester: 89, 92, and 87. How much will he have to make on the fourth exam to have a final average of an A (a 92% or better)? 100% 98% or better 96% or better 90% or better Steve has these scores on the exams for this semester: 89, 92, and 87. How much will he have to make on the fourth exam to have a final average of an A (a 92% or better)? 100% 98% or better 96% or better 90% or better @Mathematics

OpenStudy (anonymous):

100

OpenStudy (anonymous):

4. grade = x 92-89=3, 92-87=5, 92-92=0, sum of them= 8, 92+8